There are ten members of the debate team: Ann, Bill, Carla, Don, Ed, Frank, Gretchen, Harry, Isabel, and Jen. If they randomly choose three co-leaders of the group, what is the probability that Ed, Ann, and Harry are chosen as the co-leaders?
Answer #1

There are 10 members

When three person will be chosen then each one will have equal probability to be chosen.

i.e. each has probability=1/10

Now,

Probability that Ed, Ann, and Harry are chosen as the co-leaders=(1/10)*(1/10)*1/10)=0.001